Yosemite Bike Corp. manufactures mountain bikes and distributes them through retail outlets in California, Oregon, and Washington. Yosemite Bike has declared the following annual dividends over a six-year period ended December 31 of each year: 20Y1, $24,250; 20Y2, $10,500; 20Y3, $112,750; 20Y4, $97,500; 20Y5, $117,500; and 20Y6, $132,500. During the entire period, the outstanding stock of the company was composed of 25,000 shares of cumulative preferred 2% stock, $85 par, and 100,000 shares of common stock, $4 par. Required: 1. Determine the total dividends and the per-share dividends declared on each class of stock for each of the six years. There were no dividends in arrears at the beginning of 20Y1. Summarize the data in tabular form. If required, round your answers to two decimal places. If the amount is zero, please enter "0". Preferred Dividends Common Dividends Year Total Dividends Total Per Share Total Per Share 20Y1 $24,250 $ 24,250 $ 0.97 $ 0 $ 0 20Y2 10,500 10,500 0.42 0 0 20Y3 112,750 50,250 2.01 62,500 0.63 20Y4 97,500 42,500 1.7 55,000 0.55 20Y5 117,500 42,500 1.7 75,000 0.75 20Y6 132,500 42,500 1.7 90,000 0.9 $ 8.50 $ 2.83 2. Determine the average annual dividend per share for each class of stock for the six-year period. Round your answers to two decimal places. Average annual dividend for preferred stock $ 1.42 per share Average annual dividend for common stock $ 0.47 per share 3. Assuming a market price of $100 for the preferred stock and $5 for the common stock, determine the average annual percentage return on initial shareholders’ investment, based on the average annual dividend per share (A) for preferred stock and (B) for common stock. Round your answers to two decimal places. Preferred stock 0.01 % Common stock
Answer #1

1)In case of Cumulative preference shares , any preferred dividend not paid in any year is carried to future year and will be paid in preference to current year preferred and common stock dividend

Preferred dividend =par value *dividend rate

                  = (25000*85) *2%

                    = 42500

PREFERRED STOCK COMMON STOCK Preferred dividend in arrear
Total dividend paid Total per share Total per share
20Y1 24250 24250 24250/25000=.97 0 0 42500-24250= 18250
20Y2 10500 10500 10500/25000= .42 0 0 42500-10500= 32000+18250 last year = 50250
20Y3 112750 42500+50250= 92750 92750/25000= 3.71 112750-92750= 20000 20000/100000=$.20 0
20Y4 97500 42500 42500/25000 = $ 1.7 97500-42500= 55000 55000/100000=.55 0
20Y5 117500 42500 1.7 75000 75000/100000=.75 0
20Y6 132500 42500 1.7 90000 90000/100000=.90 0
Total 495000 255000 10.2 240000 2.40

2)

average annual dividend for preferred stock = Total dividend per share over 6 years /number of years

                                            = 10.2 / 6

                                             = $ 1.7 per share

Average annual dividend for common stock = 2.40 /6

                                                      = $ .40 per share

3)average annual percentage return = Dividend /market price

Preferred stock = 1.7 /100

                        = 1.7%

common stock = .40 /5

                   = .08 or 8%